# Temporal Analysis and Repair of Flaky Dockerfiles

### v1

@@ III-E RQ2: Taxonomy of Dockerfile Flakiness
Figure 1 illustrates our hierarchically structured taxonomy of Dockerfile flakiness. The left side shows the main categories of flakiness, while the right side lists the associated subcategories. The numbers within each box indicate the frequency of occurrences. The rest of this section provides an overview of the primary categories of flakiness identified in our taxonomy. Detailed information, including examples for each category and sub-category, can be found in our replication package [15].
@@
Availability and Retrieval Errors: These errors happen when previously available dependencies are not found or cannot be accessed (e.g. Retrieval/Import Issues), or decompressed due to Extraction & Invalid Format Issues. As an example of Base Image Availability Issues, one of the Docker projects we studied, Mistserver [^5] https://github.com/R0GGER/mistserver/blob/0fe477e4fb35755ad0852d46c91ed42e5b18e990/Dockerfile, uses FROM phusion/baseimage:master, leading to a failure because this specific version of the image is currently unavailable.
@@
Post-Installation Errors: After installation, flakiness can arise from the dependencies themselves, such as Vulnerability Issues, Compilation/Syntax Issues, or Runtime Issues. Additionally, improper installations can lead to Undefined Command Issues in the Dockerfile, resulting in build failures. As an example, we have observed flakiness in a Dockerfile using Symphony—a PHP framework—which is known to have vulnerability issues within its CodeExtension filters  [^6] https://github.com/symfony/symfony/security/advisories/GHSA-q847-2q57-wmr3, and using those filters may cause flakiness.
@@
Server Connectivity Errors (CON). This category is the second most prevalent among all the categories comprising 20.12% of the errors. These errors occur when there are issues while connecting to previously stable external servers. Server Access Issues arise when the Dockerfile cannot reach the target server due to invalid URLs or server downtimes. Timeout Issues occur when connections to servers take too long to establish or complete, often caused by overloaded servers or the massive size of transferred data. Internal Web Server Issues refer to errors within the accessed server, typically denoted with HTTP status codes 500s. DNS Configuration Issues occur when the Dockerfile cannot resolve the server’s domain name, leading to failed connections.
@@
Security and Authentication Errors (SEC). This category contains errors related to changes or deprecation of previous security protocols and authentication processes, making up 5.24% of the total errors. Access Control Issues arise when the Dockerfile does not have the necessary permissions to access required internal resources caused by the base image’s internal changes. Authentication/Authorization Issues manifest when there are problems verifying the identity of the user or service, which can result from incorrect or expired credentials or misconfigured authentication/authorization services. SSL/TLS Issues encompass a range of problems related to the secure transmission of data, including protocol mismatches and outdated cryptographic algorithms. GPG Key Issues arise when there are problems with the cryptographic keys used to verify the integrity and authenticity of downloaded items, which can prevent the successful retrieval and installation of necessary dependencies. Lastly, Licence Issues occur when the Dockerfile attempts to use software with new licensing restrictions.
@@
Package Manager-Related Errors (PMG). Package manager-related errors constitute 5% of the flakiness instances and refer to the changes applied to the package manager configuration during the build process. The most common subcategory of this class is Internal/Cache Issues with 4.2% of total flakiness. These errors arise from inconsistency or unreliability in the package manager’s internal system during its installation or utilization, resulting in failed operations. As an example, the command: RUN npm install --registry=r where r is no longer a reliable registry for npm would cause an internal issue within the package manager. Another subcategory is permission-denied issues, which occur when the Dockerfile does not have the necessary permissions to interact with the package manager to install or update packages. This sort of error can happen due to permission changes within the base images or other infrastructures within the Dockerfile. Lastly, Undefined package Manager Issues encompass errors caused by improper installation of package managers, resulting in a corrupted package manager within the system.
@@
Environment Errors (ENV). Environment Management Issues and Environment Configuration Issues fall into this category, representing 3.7% of the errors corresponding to interactions with virtual environments. These environment errors often arise from changes made to the base images or other underlying infrastructures specified in the Dockerfile. Such changes enforce developers to strictly adhere to the new rules to minimize vulnerabilities and enhance the system’s robustness. A detailed example of this type of error is illustrated in Listing 2 where an externally managed environment is required to alleviate the risk of disrupting the OS package management system.
@@ IV FlakiDock
Figure 2: An Overview of our proposed approach: FlakiDock
@@ V-A Implementation
FlakiDock is developed in Python. For running our experiments, we use GPT-4 model gpt-4-0613 as our LLM. For all experiments, we set the temperature parameter to 0 to ensure deterministic and well-defined answers from the LLM. Chroma DB, which is an open-source embedding database [^7] https://www.trychroma.com/, serves our need for vector similarity search. For embedding generation, we use text-embedding-ada-002 from OpenAI [31]. The temporal analysis, including project checkouts and Dockerfile builds, is performed on an infrastructure consisting of 4 Intel(R) Xeon(R) CPU 2.50GHz machines with 62 GB RAM each. For repair, we use 8 AWS machines of type t2.2xlarge, each with 8 CPUs and 32 GB RAM.
@@ V-C Baselines
We compare FlakiDock with Parfum [4] and LLM-only prompting to repair Dockerfile flakiness. We chose Parfum as it is a recent work offering automated repairs for Dockerfile smells. In contrast, other techniques, such as Shipwright [6], Hadolint [9], and Binnacle [2] focus on detecting error patterns or smells and require manual intervention for their operation. For GPT-4 Dockerfile Only, we invoke the GPT-4 model to generate repairs based solely on the Dockerfile content. Following that, we include build output to provide more context for the LLM to generate repair.
@@ V-D Results
Table II shows repair accuracy for each method across various flakiness categories. Overall, FlakiDock (With Feedback Loop) achieves the highest success rate of 73.55%, underscoring the advantage of iterative refinement in enhancing repair accuracy. Parfum exhibits the lowest accuracy of 0.58%, indicating its limited ability to address complex errors. GPT-4 Dockerfile Only slightly improves this with a 6.10% success rate, while GPT-4 Dockerfile & Build Output significantly increases effectiveness to 37.79%, highlighting the value of additional context from build outputs.
Next, we present the results per error category.
@@
Dependency-Related Errors (DEP). The repair accuracy for DEP errors with Parfum is 0% because it fails to handle the complex and dynamic nature of dependencies effectively. It relies on predefined rules that do not accommodate the diverse issues that arise from dependency changes, such as version mismatches and missing libraries. On the other hand, GPT-4 Dockerfile Only achieves a 4.64% accuracy. GPT-4 Dockerfile & Build Output increases the success rate to 37.50%, demonstrating the value of incorporating build output information to provide context for the LLM to identify and resolve dependency issues more effectively. Without a feedback loop, FlakiDock achieves a 67.86% repair accuracy, significantly outperforming the other methods by utilizing both static and dynamic information. With the inclusion of the feedback loop, FlakiDock achieves the highest accuracy of 77.14%, highlighting the advantage of iterative refinement.
Server Connectivity Errors (CON). Parfum manages a success rate of 11.11% for this error category. GPT-4 Dockerfile Only achieves a 22.22% success rate, providing moderate improvement over Parfum by utilizing LLMs to interpret server connectivity issues directly from the Dockerfile content. GPT-4 Dockerfile & Build Output improves the success rate to 44.44%, as the inclusion of build output provides additional context, allowing the LLM to understand the specific nature of server connectivity errors more effectively. FlakiDock without the feedback loop maintains the success rate at 22.22%, suggesting that the iterative feedback process is crucial for improving performance in this category. Finally, with the feedback loop, FlakiDock achieves a 44.44% success rate, comparable to the best LLM-based method.
Security & Authentication Errors (SEC). Parfum is unable to repair any issues as shown by a 0% repair accuracy. This highlights its inability to tackle the complexity of security and authentication problems, which often require more nuanced approaches than static rules can provide. GPT-4 Dockerfile only achieves a 6.25% success rate, showing a slight improvement by applying general LLM capabilities to security-related fixes, although its effectiveness is still limited due to the lack of detailed context. GPT-4 Dockerfile & Build Output sees a marginal increase in success to 12.50%, indicating some benefit from using additional build information to understand the security issues better. Furthermore, without a feedback loop, FlakiDock remains at a 6.25% success rate, indicating that the iterative feedback process is essential for tackling these complex issues effectively. With the feedback loop incorporated, FlakiDock achieves the highest repair accuracy of 37.50%.
@@
Package Manager-Related Errors (PMG). Parfum achieves a success rate of 6.25%, indicating a limited capacity to handle these flakiness category. GPT-4 Dockerfile Only performs better, with a 25% success rate. This improvement is due to the LLM’s general knowledge of package management, which enables it to address some common package-related issues based on Dockerfile content. GPT-4 Dockerfile & Build Output further improves the success rate to 37.50% by leveraging build output data to provide a clearer understanding of package manager-related problems. FlakiDock without a feedback loop maintains this 37.50% success rate. However, when the feedback loop is incorporated, the highest accuracy achieved is 56.25
@@
Environment Errors (ENV). Parfum fails to address any Environment Errors, resulting in a 0% success rate. We hypothesize that a static rule-based approach is inadequate for managing dynamic environmental issues, such as configuration changes and external environment dependencies. GPT-4 Dockerfile achieves a 4.55% success rate, reflecting a minimal ability to handle environmental factors due to its reliance solely on Dockerfile content without additional context. GPT-4 Dockerfile & Build Output significantly improves the success rate to 59.09%, benefiting from the extra context provided by build output data, which helps the LLM understand and resolve environmental errors more effectively. FlakiDock achieves an 81.82% success rate even without a feedback loop, demonstrating its strong performance in addressing environmental errors. However, incorporating a feedback loop does not provide additional improvement.
Filesystem-Related Errors (FS). For this flakiness category, all methods, including FlakiDock, fail to fix any flakiness. This consistent failure indicates that current techniques are inadequate in addressing the unique challenges posed by these errors.
VI Threats To Validity

### v2

@@ III-D RQ1: Prevalence of Dockerfile Flakiness
For instance, a project-specific error could involve a command such as RUN build_script.sh, which executes a script exhibiting non-deterministic behavior due to race conditions or reliance on external services. Similarly, RUN npm install can result in project-specific errors when dependencies listed in package.json exhibit flaky or inconsistent behavior caused by misconfigurations, version mismatches, or outdated packages. In such cases, the Dockerfile itself is correct, but the failure stems from issues within the external dependency definitions.
@@ III-E RQ2: Taxonomy of Dockerfile Flakiness
Figure 1 illustrates our hierarchically structured taxonomy of Dockerfile flakiness. The left side shows the main categories of flakiness, while the right side lists the associated subcategories. The numbers within each box indicate the frequency of occurrences. For each subcategory, we indicate its nature—deterministic, non-deterministic, or mixed—where “mixed” refers to cases that exhibit both deterministic and non-deterministic behaviors. The rest of this section provides an overview of the primary categories of flakiness identified in our taxonomy. Detailed information, including examples for each category and sub-category, can be found in our replication package [20].
@@
Availability and Retrieval Errors: These errors happen when previously available dependencies are not found or cannot be accessed (e.g. Retrieval/Import Issues) temporarily or permanently, or decompressed due to Extraction & Invalid Format Issues. As an example of Base Image Availability Issues, one of the Docker projects we studied, Mistserver [21] , uses FROM phusion/baseimage:master, leading to a failure because this specific version of the image is currently unavailable.
@@
Post-Installation Errors: After installation, flakiness can arise from the dependencies themselves, such as Vulnerability Issues, Compilation/Syntax Issues, or Runtime Issues. Additionally, improper installations can lead to Undefined Command Issues in the Dockerfile, resulting in build failures. As an example, we have observed flakiness in a Dockerfile using Symphony—a PHP framework—which is known to have vulnerability issues within its CodeExtension filters [22], and using those filters may cause flakiness.
@@
Server Connectivity Errors (CON). This category is the second most prevalent among all the categories comprising 20.12% of the errors. These errors occur when there are issues while connecting to previously stable external servers. These errors are typically non-deterministic, except when servers are permanently out of service or entirely unreachable. Server Access Issues arise when the Dockerfile cannot reach the target server due to invalid URLs or temporary server downtimes. Timeout Issues occur when connections to servers take too long to establish or complete, often caused by overloaded servers or the massive size of transferred data. Internal Web Server Issues refer to errors within the accessed server, typically denoted with HTTP status codes 500s. DNS Configuration Issues arise when the Dockerfile temporarily fails to resolve the server’s domain name, resulting in failed connections.
@@
Security and Authentication Errors (SEC). This category contains errors related to changes or deprecation of previous security protocols and authentication processes, making up 5.24% of the total errors. Access Control Issues arise when the Dockerfile does not have the necessary permissions to access required internal resources caused by the base image’s internal changes. Authentication/Authorization Issues manifest when there are problems verifying the identity of the user or service, which can result from incorrect or expired credentials or misconfigured authentication/authorization services.
This subcategory is mixed, as such errors can be affected by transient factors like token expiration or permanent external service instability.
SSL/TLS Issues encompass a range of problems related to the secure transmission of data, including protocol mismatches and outdated cryptographic algorithms. GPG Key Issues arise when there are problems with the cryptographic keys used to verify the integrity and authenticity of downloaded items, which can prevent the successful retrieval and installation of necessary dependencies. The nature of these errors is mixed, as they can result from invalid keys or temporarily unavailable key servers. Lastly, Licence Issues occur when the Dockerfile attempts to use software with new licensing restrictions.
@@
Package Manager-Related Errors (PMG). Package manager-related errors constitute 5% of the flakiness instances and refer to the changes applied to the package manager configuration during the build process. The most common subcategory of this class is Internal/Cache Issues with 4.2% of total flakiness. These errors arise from inconsistency or unreliability in the package manager’s internal system during its installation or utilization, resulting in failed operations. This subcategory is mixed, as deterministic failures can stem from corrupted caches, while non-deterministic failures may result from transient issues with the package manager’s internal operations. As an example, the command: RUN npm install --registry=r where r is no longer a reliable registry for npm would cause an internal issue within the package manager. Another subcategory is permission-denied issues, which occur when the Dockerfile does not have the necessary permissions to interact with the package manager to install or update packages. This sort of error can happen due to permission changes within the base images or other infrastructures within the Dockerfile. Lastly, Undefined package Manager Issues encompass errors caused by improper installation of package managers, resulting in a corrupted package manager within the system.
@@
Environment Errors (ENV). Environment Management Issues and Environment Configuration Issues fall into this category, representing 3.7% of the errors corresponding to interactions with virtual environments. These environment errors often arise from changes made to the base images or other underlying infrastructures specified in the Dockerfile. Such changes enforce developers to strictly adhere to the new rules to minimize vulnerabilities and enhance the system’s robustness. A detailed example of this type of error is illustrated in Listing 1 where an externally managed environment is required to alleviate the risk of disrupting the OS package management system.
@@ IV FlakiDock
Figure 2: Overview of our proposed approach: FlakiDock
@@ V-A Implementation
FlakiDock is developed in Python. For our experiments, we use the GPT-4 model (gpt-4-0613) as the LLM, with the temperature parameter set to 0 to ensure deterministic and well-defined outputs. Chroma DB, which is an open-source embedding database [39], serves our need for vector similarity search. For embedding generation, we use text-embedding-ada-002 from OpenAI [38]. The temporal analysis, including project checkouts and Dockerfile builds, is performed on an infrastructure consisting of 4 Intel(R) Xeon(R) CPU 2.50GHz machines with 62 GB RAM each. For repair, we use 8 AWS machines of type t2.2xlarge, each with 8 CPUs and 32 GB RAM.
@@ V-C Baselines
We compare FlakiDock with Parfum [5], Shipwright [7] and LLM-only prompting to repair Dockerfile flakiness. We chose Parfum as it is a recent work offering automated repairs for Dockerfile smells. Shipwright uses 13 repair-pattern rules to automatically repair Dockerfiles that fail to build. In contrast, other techniques, such as Hadolint [10] and Binnacle [3] focus on detecting error patterns or smells and require manual intervention for their operation. For GPT-4 Dockerfile Only, we invoke the GPT-4 model to generate repairs based solely on the Dockerfile content. Following that, we include build output to provide more context for the LLM to generate repairs.
@@ V-D Results
Table II presents the repair accuracy of each method across flakiness categories. FlakiDock (With Feedback Loop) achieves the highest success rate of 73.55%, demonstrating the effectiveness of iterative refinement. In comparison, Parfum has the lowest accuracy at 0.58% and Shipwright achieves 5.52%, both highlighting their limitations in addressing complex errors. GPT-4 Dockerfile Only achieves 6.10%, while GPT-4 Dockerfile & Build Output improves significantly to 37.79%, emphasizing the importance of incorporating build output context.
@@
Dependency-Related Errors (DEP). The repair accuracy for DEP errors with Parfum is 0%, as it relies on predefined rules that fail to address the complex and dynamic nature of dependencies, such as version mismatches or missing libraries. Shipwright achieves 6.79%, which is notable as this is the only category it can handle, highlighting its limitations in addressing broader flakiness issues. GPT-4 Dockerfile Only achieves 4.64% accuracy, while GPT-4 Dockerfile & Build Output improves to 37.50% by leveraging build output context, enabling the LLM to better identify and resolve dependency issues. FlakiDock, without a feedback loop, achieves 67.86% accuracy, significantly outperforming other methods. With the inclusion of feedback loop, FlakiDock yields the highest accuracy of 77.14%, demonstrating the effectiveness of iterative refinement in resolving dependency-related errors.
Server Connectivity Errors (CON). Parfum achieves an accuracy of 11.11% for addressing server connectivity errors. GPT-4 Dockerfile Only improves this to 22.22%, while the inclusion of build output further increases the accuracy to 44.44%, leveraging additional context for a more effective resolution. FlakiDock, incorporating its feedback loop, achieves an equivalent accuracy of 44.44%, demonstrating performance comparable to the most effective LLM-based approach for this category.
Security & Authentication Errors (SEC). Security and authentication errors pose significant challenges due to their complexity, often requiring nuanced approaches beyond static rules. GPT-4 Dockerfile achieves a 6.25% success rate, reflecting limited effectiveness due to its reliance on general LLM capabilities without detailed context. Incorporating build output increases accuracy to 12.50%, highlighting the value of additional contextual information. Without a feedback loop, FlakiDock also achieves 6.25%, underscoring the importance of iterative refinement. With the feedback loop, FlakiDock attains the highest repair accuracy of 37.50%, demonstrating its ability to effectively address these intricate issues.
@@
Package Manager-Related Errors (PMG). Parfum achieves a modest success rate of 6.25%, reflecting its limited ability to handle this category. GPT-4 Dockerfile Only performs better, with a 25% success rate. This improvement is due to the LLM’s general knowledge of package management, which enables it to address some common package-related issues based on Dockerfile content. Incorporating build output raises the accuracy to 37.50%, as the additional context enables better understanding of package manager-related problems. FlakiDock, without a feedback loop, matches this 37.50% accuracy. However, when the feedback loop is incorporated, the highest accuracy achieved is 56.25
@@
Environment Errors (ENV). GPT-4 Dockerfile achieves a 4.55% success rate, demonstrating limited effectiveness due to its reliance solely on Dockerfile content. Incorporating build output raises the success rate to 59.09%, as the additional context allows the LLM to better understand and resolve environmental issues. FlakiDock outperforms all other methods, achieving an 81.82% success rate even without a feedback loop, demonstrating its strong performance in addressing environmental errors. However, incorporating a feedback loop does not provide additional improvement.
Filesystem-Related Errors (FS). For this flakiness category, all techniques, including FlakiDock, fail to resolve any flakiness. The primary challenge lies in the inability of existing tools and LLMs to account for changes in the updated image’s filesystem, such as a modified WORKDIR. Although feedback or similar examples could potentially aid in addressing these issues, our evaluation is limited to a single example, making it difficult to draw generalized conclusions about this category.
VI Discussion
(Non-)Deterministic Dockerfile Flakiness. Dockerfile flakiness can manifest as deterministic or non-deterministic temporal failures, each requiring tailored repair approaches. Deterministic flakiness results in persistent and reproducible build errors caused by specific external changes, such as the deprecation of a dependency or removal of a required resource. For example, in Listing 3, a deterministic failure occurs due to the unavailability of the npm.taobao.org host, causing the wget command to fail. FlakiDock analyzes the static and dynamic build information, identifies the issue, and proposes a repair by substituting the unavailable URL with a valid one, as shown in the repaired Dockerfile. However, even after resolving a deterministic error, it is important to note that new issues may arise in the future due to the evolving nature of Dockerfile dependencies and environments.
Build Output Summary
wget: unable to resolve host address ’npm.taobao.org’
Repaired Dockerfile
- RUN wget https://npm.taobao.org/mirrors/node/v8.9.1/node-v8.9.1-linux-x64.tar.gz && \
+ RUN wget https://nodejs.org/dist/v8.9.1/node-v8.9.1-linux-x64.tar.gz && \
tar -C /usr/local --strip-components 1 -xzf node-v8.9.1-linux-x64.tar.gz && \
rm node-v8.9.1-linux-x64.tar.gz
Listing 3: Host Resolution Error (Deterministic)
In contrast, non-deterministic flakiness arises sporadically due to transient factors, such as server connectivity interruptions or temporary package manager caching inconsistencies. These failures lack a consistent pattern, complicating root-cause analysis and repair validation. For instance, Listing 4 demonstrates a timeout error caused by a transient network issue during a yarn install command. To address this, FlakiDock adjusts the network timeout configuration in the Dockerfile to ensure stability, as shown in the repaired version. While FlakiDock can propose an initial repair based on observed data, a single successful repair validation is insufficient. Instead, repairs for non-deterministic failures must be tested iteratively over time to ensure reliability and prevent reoccurrence. The repair validator in FlakiDock facilitates this by running repeated builds, monitoring outcomes across multiple iterations, and confirming the robustness of the repair.
Build Output Summary
An unexpected error occurred: "https://registry.yarnpkg.com/@material-ui/icons/-/icons-4.11.3.tgz: ESOCKETTIMEDOUT".
Repaired Dockerfile
FROM node:16.14.0
COPY --from=deps /tmp/deps.json ./package.json
COPY yarn.lock ./
+ RUN yarn config set network-timeout 600000
RUN yarn install --production
Listing 4: Timeout Error (Non-Deterministic)
In FlakiDock, deterministic flakiness is repaired with greater certainty due to its consistent root cause. Non-deterministic flakiness, however, requires iterative testing and extended validation to ensure repair effectiveness. FlakiDock addresses both by offering immediate repairs for deterministic flakiness and incorporating a time-based validation process for non-deterministic cases, adjusting parameters such as $n$ for iterative validation.
Threshold $n$ for Determining a Successful Repair. The parameter $n$, representing the number of build iterations used to validate repairs, is crucial for ensuring repair accuracy in addressing Dockerfile flakiness. In our study, $n$ was set to two based on empirical analysis of the demonstration dataset, capturing over 90% of flaky behaviors. For temporal deterministic errors, this value is sufficient, as these errors are resolved once the root cause is addressed. Once a deterministic error has been repaired successfully in a single build, it is unlikely to reoccur, so additional iterations have little effect on repair accuracy.
Non-deterministic flakiness presents a more complex challenge. Since these errors can occur unpredictably, passing two rounds of validation may not suffice to reliably capture and resolve the flakiness, requiring a higher $n$ to ensure robustness. A potential enhancement to FlakiDock would be the inclusion of a detector that identifies the type of Dockerfile flakiness (deterministic or non-deterministic). Based on this classification, and based on the retrieved similar records, the parameter $n$ could be adjusted accordingly. For deterministic flakiness, a lower $n$ is sufficient, while non-deterministic flakiness would require a higher $n$ to ensure a reliable repair. This would prevent premature validation of repairs for non-deterministic errors.
Dynamic Nature of Dockerfile Flakiness. Dockerfiles rely on evolving components such as base images, package managers, libraries, external URLs, and infrastructure. Changes in these dependencies such as deprecated libraries, broken URLs, or updated environment variables, can introduce flakiness, causing builds to fail. Conversely, flakiness can be resolved as dependencies stabilize without Dockerfile modifications, highlighting the temporal and dynamic nature of Dockerfile flakiness. Our nine-month study revealed the temporal nature of Dockerfile flakiness, where initially stable Dockerfiles became flaky due to dependency updates, and flaky ones stabilized as external issues were resolved. Static analysis tools fail to detect such failures, as they lack awareness of dynamic dependency changes (e.g., Listings 1 and 2). We propose a taxonomy to help developers systematically diagnose and address flakiness while preparing for future instability. This underscores the need for FlakiDock, which dynamically adapts to evolving dependencies by providing context-aware repairs.
Limitations of Existing Tools for Dockerfile Flakiness. Existing tools for Dockerfile repair, such as Shipwright and Parfum, exhibit significant limitations when addressing the dynamic and multifaceted nature of Dockerfile flakiness. Shipwright achieved only 19 repairs (5.52%) in our evaluation. However, this represents an upper bound, as a repair is deemed successful if Shipwright generates an output Dockerfile using its predefined patterns, irrespective of whether it fully addresses the underlying flakiness. This tool relies on 13 repair patterns, heuristically designed by its authors through a clustering approach. These patterns must be maintained and updated to remain effective as new failure types emerge. Without regular updates, the effectiveness of the tool is likely to decline, given the dynamic and evolving nature of Dockerfile dependencies, which often introduce novel and unforeseen failure scenarios. Shipwright was capable of handling only Dependency-Related Errors (DEP), likely because such errors often align with its predefined patterns, such as missing or incorrect dependencies. Similarly, Parfum focuses on repairing Dockerfile smells but lacks mechanisms to address dynamic flakiness, achieving only a 0.58% repair accuracy. In contrast, our proposed tool, FlakiDock, integrates both static and runtime contexts using retrieval-augmented generation and an iterative feedback loop, achieving a significantly higher repair accuracy of 73.55%. These results highlight the shortcomings of static, heuristic-driven tools and emphasize the need for dynamic, context-aware approaches to effectively handle the complexities of Dockerfile flakiness. However, FlakiDock is not without limitations. It relies on LLMs, which may struggle with uncommon flakiness issues that fall outside their training scope. Its effectiveness depends on the diversity and quality of the demonstration dataset, limiting performance in addressing rare error categories. Additionally, the iterative repair process, while effective, can be computationally intensive and time-consuming.
VII Threats To Validity